# Model-independent determination of nuclear charge radii from Li-like ions

### v1

@@ I Introduction
The nuclear root-mean-square (rms) charge radius is a fundamental parameter that determines the deviation of the electrostatic interaction between the nucleus and atomic electrons from its pure Coulomb form. Its accurate knowledge is important for tests of bound-state QED [1], understanding of nuclear beta-decay [2, 3], and for studies of parity violation in atomic systems [4]. Since accurate ab initio calculations of nuclear radii are currently not feasible, the values available in the literature are determined from experiment, primarily from electron–nucleus scattering [5] and spectroscopy of muonic atoms [6].
@@ V Nuclear radii determination from Li-like transition energies
We now turn to the determination of the nuclear charge radii from the $2p_{j}\,$- 2s\,2s transition energies of Li-like ions. High-precision theoretical predictions for these transition energies were recently reported in Ref. [23] for the whole isoelectronic sequence with $Z\geq 10$. Accurate experimental results are also available, but mostly for low- and medium-$Z$ ions. In the high-$Z$ region, 208Pb and 209Bi are the only elements for which both electron–scattering data and precisely measured transition energies are presently available. In the following, we determine the nuclear charge radii for these two elements.
@@ V.1 Charge radius of 208Pb
In order to determine $r_{C}$, we calculate the field-shift differences $\delta_{\rm fs}E$ numerically by slightly varying the parameters of our nuclear model around the initial point (7) and recording the values of $\delta_{\rm fs}E$, $r_{C}$, $V_{24}$, and $V_{26}$ in each case. In our calculations of $\delta_{\rm fs}E$, we included the Dirac energies, the one-photon exchange correction, and the dominant parts of the two-photon exchange and one-loop QED effects; see Ref. [23] for details. We then fit the numerical results for $\delta_{\rm fs}E$ to the form [43]
$\delta_{\rm fs}E=C_{2}\,\delta\langle r^{2}\rangle+C_{4}\delta\langle r^{4}\rangle+C_{6}\delta\langle r^{6}\rangle\,$, (10)
where $\delta\langle r^{2n}\rangle=\langle r^{2n}\rangle-\langle r^{2n}\rangle_{0}$ . We note that the above representation is not a proper expansion, since it does not correspond to the analytical form of the fns correction of Eq. (II). We use the ansatz (10) solely to interpolate our numerical data, with the interpolation error remaining negligible as long as the interpolation region is kept sufficiently small.
In practical calculations, we computed a set of $\delta_{\mathrm{fs}}E$ values with the 3pF nuclear model, varying the nuclear parameters by typically 0.010.01-$0.02$ fm, and then fitted our results to the form (10). The analysis performed in Sec. III shows that the actual choice of the nuclear model is irrelevant in this context. We found that in this (very small) parameter region, the interpolation formula (10) yields an essentially exact fit to our numerical results (though its precision decreases rapidly as the parameter range is expanded).
Specifically, our fit yields the coefficients $C_{2}=-0.288$ eV/fm2, $C_{4}=3.25\times 10^{-4}$ eV/fm4, and $C_{6}=-6.5\times 10^{-7}$ eV/fm6 for the $D_{1}$ transition, as well as $C_{2}=-0.311$ eV/fm2, $C_{4}=3.48\times 10^{-4}$ eV/fm4, and $C_{6}=-6.9\times 10^{-7}$ eV/fm6 for the $D_{2}$ transition. We note that the ratios of these coefficients $C_{4}/C_{2}$ and $C_{6}/C_{2}$ are close to the values previously reported by Seltzer [43] and Blundell et al. [44].
@@ V.2 Charge radius of 209Bi
By calculating the field-shift differences $\delta_{\mathrm{fs}}E$ and fitting the numerical results to the form (10), we obtain $C_{2}=-0.3442$ eV/fm2, $C_{4}=3.91\times 10^{-4}$ eV/fm4, and $C_{6}=-7.7\times 10^{-7}$ eV/fm6. With the help of these coefficients, we immediately obtain our result for the rms radius of 209Bi,

### v2

@@ I Introduction
The nuclear root-mean-square (rms) charge radius is a fundamental parameter that determines the deviation of the electrostatic interaction between the nucleus and atomic electrons from its pure Coulomb form. Its accurate knowledge is important for tests of bound-state QED [1], understanding of nuclear beta-decay [2, 3], determining the nuclear neutron-distribution radii and properties of neutron-star crusts [4, 5], as well as for studies of parity violation in atomic systems [6]. Since accurate ab initio calculations of nuclear radii are currently not feasible, the values available in the literature are determined from experiment, primarily from electron–nucleus scattering [7] and spectroscopy of muonic atoms [8].
@@ V Nuclear radii determination from Li-like transition energies
We now turn to the determination of nuclear charge radii from the $2p_{j}\,$ 2s\,2s transition energies in Li-like ions. High-precision theoretical predictions for these transitions were recently reported in Ref. [25] for the entire isoelectronic sequence with $Z\geq 10$. Theoretical energies become increasingly sensitive to the nuclear radii as the nuclear charge $Z$ increases, thus making high-$Z$ ions promising candidates for a competitive determination of nuclear radii. Among the heavy elements, 208Pb and 209Bi are currently the only cases for which both electron-scattering data and precise spectroscopic measurements of the relevant transition energies are available. In the following, we extract the nuclear charge radii for these two nuclei.
@@ V.1 Charge radius of 208Pb
In order to determine $r_{C}$, we calculate the field-shift differences $\delta_{\rm fs}E$ numerically by slightly varying the parameters of our nuclear model around the initial point (7) and recording the values of $\delta_{\rm fs}E$, $r_{C}$, $V_{24}$, and $V_{26}$ in each case. We then fit the numerical results for $\delta_{\rm fs}E$ to the form
$\delta_{\rm fs}E=C_{2}\,\delta\langle r^{2}\rangle^{\gamma}+C_{4}\,\delta\langle r^{4}\rangle+D_{4}\,\delta\langle r^{2}\rangle^{2}+C_{6}\,\delta\langle r^{6}\rangle\,$, (10)
where $\delta\langle r^{2n}\rangle=\langle r^{2n}\rangle-\langle r^{2n}\rangle_{0}$, $\gamma=\sqrt{1-({Z\alpha})^{2}}$, and the coefficients $C_{n}$ and $D_{4}$ are free parameters determined by fitting. We note that the ansatz (10) differs from the one used by Seltzer [45] in that we included $\gamma$ in the $C_{2}$ part and added the $D_{4}$ contribution. This ansatz complies better with the analytical formula (II) and the high-$Z$ $r_{C}^{2\gamma}$ scaling. As a consequence, it yields a better fit for our numerical data. Nevertheless, we emphasize that Eq. (10) is not a rigorous analytical expansion but merely an approximate representation. As a consequence, the coefficients in Eq. (10) slightly depend on the reference point (7) and the chosen fitting range. We employ Eq. (10) only for interpolation of our numerical data and have verified that the interpolation errors are fully negligible for our purposes.
In practical calculations, we computed a set of $\delta_{\mathrm{fs}}E$ values with the 3pF nuclear model, and then fitted our results to the form (10). The analysis performed in Sec. III shows that the actual choice of the nuclear model is irrelevant in this context. In our calculations of $\delta_{\rm fs}E$, we included the Dirac energies, the one-photon exchange correction, as well as the dominant parts of the two-photon exchange and one-loop QED effects; see Ref. [25] for details.
Specifically, our fit yields the coefficients $C_{2}=-0.662$ eV/fm2, $C_{4}=3.22\times 10^{-4}$ eV/fm4, $D_{4}=-3.1\times 10^{-4}$ eV/fm4, and $C_{6}=-6.4\times 10^{-7}$ eV/fm6 for the $D_{1}$ transition, as well as $C_{2}=-0.715$ eV/fm2, $C_{4}=3.47\times 10^{-4}$ eV/fm4, $D_{4}=-3.4\times 10^{-4}$ eV/fm4, and $C_{6}=-6.9\times 10^{-7}$ eV/fm6 for the $D_{2}$ transition.
@@ V.2 Charge radius of 209Bi
By calculating the field-shift differences $\delta_{\mathrm{fs}}E$ and fitting the numerical results to the form (10), we obtain $C_{2}=-0.810$ eV/fm2, $C_{4}=3.90\times 10^{-4}$ eV/fm4, $D_{4}=-3.9\times 10^{-4}$ eV/fm4, and $C_{6}=-7.6\times 10^{-7}$ eV/fm6. With the help of these coefficients, we obtain our result for the rms radius of 209Bi,